Assume you add 175.0 mL of water to 125.0 mL of a 0.361 M KNO3 solution. What is the final molarity of your diluted solution?
1
Expert's answer
2019-05-09T07:21:46-0400

"n(KNO3)=C*V=0.361M*0.125L=0.045125 moles"

"C(KNO3)=n\/V=0.045125 moles\/0.125L+0.175L=0.15mol\/L"